OpenStudy (anonymous):

Find the average rate of change of the function f over the given interval. f(x) = 0.25x4 - x2 - 5x + 5 from x = - 9 to x = 5

OpenStudy (anonymous):

the average change in f can be found by [f(5) - f(-9)] / (5-(-9))

OpenStudy (anonymous):

i'm getting -115.23, but its not right and i'm not sure what i'm doing wrong.

OpenStudy (anonymous):

Find The Derivative: -5-2 x+x^3 Solve For x = 5 -5-2*5+5^3=110 [1] Solve For x = -9 -5-2*-9-9^3=-716 [2] [1]-[2]= 110--716=826 I think that should work its the Definite Derivative :)

OpenStudy (anonymous):

it didnt work. :(

OpenStudy (anonymous):

LOL sorry yes i made a mistake Find The Derivative: -5-2 x+x^3 This represents the instantaneous rate of change for any value of x But you need to do is calculate the gradient of the line between the tow points

OpenStudy (anonymous):

http://home.windstream.net/okrebs/page201.html This Should Help